OpenStudy (anonymous):

The sum of two consecutive integers is 141. Find the integers

OpenStudy (texaschic101):

do you know what consecutive means ?

OpenStudy (texaschic101):

Consecutive means one right after the other. So lets make 1 number be x, making the other number be x + 1. so we have : x + (x + 1) = 141 -- combine like terms 2x + 1 = 141 -- subtract 1 from both sides 2x = 140 -- divide both sides by 2 x = 70 making x + 1 = 70 + 1 = 71 check.. 70 + 71 = 141 141 = 141 (correct) your 2 numbers are 70 and 71 any questions ?

OpenStudy (anonymous):

consecutive integers are a sequence of integers that increase by one.

OpenStudy (texaschic101):

sorry..I kinda jumped the gun on you....well...do you understand what I did ?